# 0000-0002-5763-8174

### v1

@@ Abstract
[Abstract] Purpose: Magnetic Resonance Fingerprinting (MRF) enables rapid quantitative imaging, but high-resolution 3D reconstructions remain computationally expensive due to repeated NUFFTs, and the commonly used Locally Low Rank (LLR) regularization becomes ineffective at high acceleration. Learned 3D priors could address these limitations, but training them at scale is challenging due to memory and runtime constraints. This work proposes SPUR-iG, a fully 3D deep unrolled subspace reconstruction framework for efficient and scalable 3D MRF.
Methods: SPUR-iG leverages implicit GROG-based data consistency (DC), which grids non-Cartesian k-space using a learned kernel, enabling efficient FFT-based DC with minimal artifacts. To enable scalable 3D unrolled training, we introduce a staged training strategy that keeps computation tractable while progressively improving reconstruction quality. We evaluate the method on a large in vivo dataset, as well as on cross-vendor out-of-distribution data.
@@
Conclusion: SPUR-iG introduces a fully 3D unrolled reconstruction framework for MRF that improves both reconstruction speed and accuracy, making high-resolution accelerated 3D MRF more practical for clinical and research use.
@@ 1 Introduction
Magnetic Resonance Fingerprinting (MRF) ma2013magnetic is a qMRI technique that encodes tissue parameters through a sequence of varying flip angles (FA), repetition times (TR), and potentially other acquisition settings across $N_{\text{TR}}$ time points. At each TR, an image is acquired using an incoherency-promoting trajectory (e.g., rotated spirals), producing artifacts that are spatially and temporally incoherent and thus reduce interference with tissue-specific signal evolutions. After reconstructing the time series of images, voxel-wise signal evolutions are matched to a dictionary of Bloch-simulated signals, linking tissue parameters to their expected dynamics and enabling parameter estimation.
@@
Subspace reconstruction zhao2018improved, asslander2018low, liang2007spatiotemporal reduces the computational load by approximating the signal dictionary with a low-rank subspace of size $k\ll N_{\text{TR}}$, thereby lowering dimensionality and implicitly regularizing the reconstruction. The resulting subspace coefficient maps provide a compact representation of the image’s signal evolutions across TRs, which can be used to extract quantitative $T_{1}$ and $T_{2}$ maps. Beyond traditional parameter mapping, these coefficient maps capture rich signal information that facilitates additional downstream applications. These range from multi-compartment modeling nagtegaal2020fast, which is particularly valuable at gray–white matter (GM/WM) boundaries tang2018multicompartment, to high-quality clinical contrast synthesis yurt2024unlocking, wang2023high by capturing signal components not typically described by Bloch-simulated dictionaries (e.g., magnetization transfer). These advantages motivate a focus on reconstructing accurate subspace coefficient maps that can serve as input for diverse applications beyond conventional single-compartment parameter mapping.
Yet, for highly accelerated MRF, additional regularization beyond the subspace model is required. Locally Low Rank (LLR) tamir2017t2, lima2019sparsity is widely used, enforcing low-rank structure within local patches, but it requires computationally expensive singular value decompositions (SVD) cai2010singular. At even higher acceleration rates, such as acquisitions targeting 1 mm isotropic whole-brain mapping in under 2 minutes, LLR alone is insufficient to preserve reconstruction quality.
Broadly, achieving high-fidelity 3D MRF at this resolution and these acceleration factors is limited by two primary bottlenecks: (1) the reliance on handcrafted priors that lack the expressivity to resolve fine details under high undersampling, and (2) the computationally expensive NUFFTs and regularization operators, which substantially raise reconstruction time and memory usage at high resolutions.
@@
However, even with iGROG, reconstruction still relies on suboptimal LLR priors. To address this, state-of-the-art unrolled methods like SOTIP zou2025improved utilize deep learning to incorporate learned priors, showing significant improvements in both quality and speed. These methods interleave 3D data-consistency (DC) steps with denoising; however, due to memory constraints, the denoising is applied slice-by-slice in 2D. Hence, this approach fails to exploit the inherently 3D nature of MRF acquisitions, which exhibit 3D artifacts and spatial correlations. Furthermore, to accelerate reconstruction, SOTIP employs NUFFTs with reduced kernel sizes and oversampling, introducing small artifacts that the denoiser must learn to remove.
To address these issues, we propose SPUR-iG (Staged Pretraining for Unrolled Reconstruction with iGROG), a fully 3D deep unrolled reconstruction framework aggarwal2018modl, adler2018learned. Our approach combines efficient iGROG-based DC with a learned 3D prior and a staged training strategy that enables large-scale unrolled training. We first pretrain a standalone denoiser with extensive data augmentation, then train with a greedy per-unroll-iteration loss to reduce memory demands ozturkler2022gleam, and finally fine-tune the full unrolled model using gradient checkpointing chen2016training. To further improve robustness, the denoiser takes as input both the image and the unroll iteration index, enabling it to amortize performance across varying artifact and noise levels. Evaluated on 45 in vivo acquisitions with retrospective undersampling spanning acquisition times from 2 minutes down to 30 seconds, and including an example from an out-of-distribution scanner, our framework achieves whole-brain 1 mm isotropic reconstructions in under 15 s in all cases, while surpassing existing methods in quality.
@@ 2 Methods
To enable fast and accurate MRF reconstruction, we build on the standard subspace formulation and address its limitations of suboptimal priors and long reconstruction times. We propose a fully 3D unrolled framework that integrates a learned prior with an efficient DC formulation, along with a staged training strategy that makes large-scale training feasible within a reasonable compute and time budget. We begin by reviewing MRF acquisition and the subspace model, followed by unrolled reconstruction, and then present the details of our proposed method and evaluation pipeline. An overview of our approach is illustrated in Fig. 1.
Figure 1:  Illustration of inference (top) and training (bottom) pipelines of SPUR-iG. During inference, iGROG grids the non-Cartesian data onto a Cartesian grid using an implicit kernel representation learned from the calibration region. The gridded k-space is then passed to the unrolled reconstruction, initialized with ${\bf\cal A}^{H}{\bf b}$. Reconstruction proceeds for NN unroll iterations, each alternating between a DC update with a learnable step size $\mu_{i}$ and a step-conditioned denoiser. Training is performed in three stages: (1) a step-conditioned UNet is pretrained on inputs with varying artifact levels, using a heuristic mapping between input type and unroll step to promote robustness across noise/artifact conditions; (2) the pretrained UNet initializes GLEAM-style training, where losses are computed and backpropagated after each unroll iteration, detaching subsequent steps; and (3) full unrolled training, initialized from stage (2), fine-tunes the model under the exact inference setting, with gradient checkpointing (GC) across DC and denoising steps to reduce memory consumption.
@@ 2.1 MRF Subspace Reconstruction
MRF consists of a sequence of $N_{\text{TR}}$ acquisitions with varying excitation parameters (e.g., FAs), producing a time series of images ${\bf x}\in\mathbb{C}^{N_{\text{TR}}\times N_{v}}$, where $N_{v}$ denotes the number of voxels in the volume. The excitation schedule is typically designed to maximize separability between tissue parameter pairs $(T_{1},T_{2})$, and reconstruction proceeds in two stages: (1) recovery of the time series $\hat{{\bf x}}$ from undersampled k-space data, and (2) voxel-wise parameter estimation via dictionary matching, where the temporal signal evolution is compared to a precomputed dictionary, and quantitative maps are obtained by selecting the entry with maximum cosine similarity.
Directly reconstructing the full time series $\bf x$} is computationally intensive and ill-posed due to its high dimensionality and the heavy undersampling used in MRF. This can be mitigated by exploiting temporal redundancy and projecting the data onto a low-dimensional subspace ${\bm{\Phi}}\in\mathbb{C}^{N_{\text{TR}}\times k}$ ($k\ll N_{\text{TR}}$) spanned by the top $k$ singular vectors of the MRF dictionary. The time series is then approximated as ${\bf x}\approx{\bm{\Phi}}{\bm{\alpha}}$, where ${\bm{\alpha}}\in\mathbb{C}^{k\times N_{v}}$ are the subspace coefficient maps. We define the forward operator acting on $\bm{\alpha}$} as ${\bf\cal A}={\bf\cal F}_{u}{\bf S}{\bm{\Phi}}$, with $\bf S$} the coil sensitivities and ${\bf\cal F}_{u}$ the undersampled Fourier operator. Since ${\bf\cal F}_{u}{\bf S}$ acts on the spatial domain and $\bm{\Phi}$} on the temporal domain, these operators commute, allowing the costly ${\bf\cal F}_{u}{\bf S}$ to be applied directly in the reduced $k$-dimensional space.
@@ 2.2 Unrolled Reconstruction
Equation (1) is commonly solved using accelerated proximal algorithms such as FISTA beck2009fast, which alternate between DC updates and proximal steps. However, at high acceleration rates, LLR regularization becomes insufficient to constrain the undersampled problem, and in addition, its proximal operator requires repeated SVDs of local patches, making it computationally expensive.
Unrolled reconstruction addresses these issues by replacing the proximal operator with a learned denoiser, yielding the update scheme
@@ 2.3.1 Efficient Data Consistency via iGROG
The main cost in (2) arises from evaluating ${\bf\cal A}^{H}{\bf\cal A}(\cdot)$, which relies on NUFFTs. To accelerate this step, we use iGROG, which learns an INR of the gridding kernel. Using calibration data, the INR is trained to map $n$ offset vectors ${\bf d}_{1},\ldots,{\bf d}_{n}$, corresponding to displacements between non-Cartesian samples and a target Cartesian grid point, to a kernel $G_{{\bm{\psi}}^{*}}({\bf d}_{1},\ldots,{\bf d}_{n})$. For each target grid location, nearby source samples are selected, the kernel is evaluated at their relative displacements, and the resulting weights are used to compute the target k-space value. This enables GRAPPA-like griswold2002generalized interpolation and, through the use of an INR, generalizes across the diverse orientations encountered in the trajectory. This allows us to first grid the non-Cartesian data onto a Cartesian grid, after which NUFFTs can be replaced by efficient FFTs. The upper left part of Fig. 1 illustrates this process.
The approach provides flexibility to trade off accuracy and noise amplification against speed. By using an oversampled grid, interpolation quality improves (since the interpolation distances are smaller), but reconstruction time increases. In our experiments, we used a $\times 1.5$ oversampling ratio with 5 input points per interpolation, achieving high-quality reconstructions while retaining substantial speed-ups. The INR training is fast (under 25 s) and can be performed during MRF acquisition immediately after calibration data is acquired. Once trained, gridding the full dataset is also efficient (under 3 s), yielding data on a Cartesian grid for subsequent FFT-based DC updates.
@@ 2.3.2 Efficient Training of a 3D Unrolled Network
To enable training of a 3D unrolled network in the challenging MRF reconstruction setting, we adopt a progressive three-stage strategy, illustrated in the lower part of Fig. 1. This strategy consists of the following stages:
@@
    Denoiser pretraining: We begin by pretraining a standalone, iteration-conditioned 3D UNet on a diverse set of inputs, where the model is trained to recover clean images from noisy, undersampled inputs. To obtain a robust initialization for subsequent stages, we expose the network to a wide range of artifact and noise levels at its input, including zero-filled (${\bf\cal A}^{H}{\bf b}$) and LLR-based reconstructions from scans with different acquisition durations (30 s, 1 min, 2 min, and 6 min, denoted ${\bm{\alpha}}_{\text{LLR}}^{\text{30s}},\ldots,{\bm{\alpha}}_{\text{LLR}}^{\text{6min}}$). Since the UNet is iteration-conditioned, but this stage does not involve actual unrolling and therefore has no natural iteration index, we assign heuristic pseudo-iteration indices to each input type. Specifically, ${\bf\cal A}^{H}{\bf b}$ is paired with index 1, corresponding to the initialization of unrolled reconstruction, while reconstructions with increasing acquisition durations are mapped to higher indices (e.g., ${\bm{\alpha}}_{\text{LLR}}^{\text{30s}}\to 2$). Because this stage does not involve unrolling or DC updates, we can efficiently apply extensive data augmentation, including random rotations, translations, and scaling. At the end of this stage, the network provides a well-initialized denoising prior that generalizes across a range of artifact levels and incorporates a notion of the unroll iteration index.
@@
    Greedy unrolled training (GLEAM): Starting from the pretrained weights, we train the unrolled network using a greedy scheme ozturkler2022gleam, in which the loss is backpropagated through only a single unroll iteration at a time, after which the computation graph is detached. This avoids storing the full unrolled graph, which would otherwise exceed typical GPU memory. Since the final output of the unrolled process is of primary interest, we assign geometrically increasing loss weights across iterations, emphasizing later steps while allowing greater flexibility in earlier ones. Importantly, this stage resolves the input distribution mismatch introduced during prior pretraining: while the denoiser from Stage 1 is trained on reconstructed images with different artifacts levels, during unrolled inference its true inputs are the intermediate outputs of preceding iterations. By the end of this stage, we obtain a fully unrolled model that achieves high reconstruction quality, with faster training convergence and improved performance due to the informed initialization.
@@
Collectively, these stages enable efficient large-scale training by progressively mitigating the limitations of each preceding stage, while circumventing the need to train a fully 3D unrolled model from scratch, which would otherwise be prohibitively compute expensive and slow. This process ultimately yields a high-quality, fully 3D unrolled reconstruction model. We refer to the resulting framework as SPUR-iG, introduced in Section 1. Code and trained models will be made publicly available upon publication.

### v2

@@ Abstract
[Abstract] Purpose: Magnetic Resonance Fingerprinting (MRF) enables rapid quantitative imaging, but high-resolution 3D reconstructions remain computationally expensive due to repeated NUFFTs, and the commonly used Locally Low Rank (LLR) regularization becomes ineffective at high acceleration. Learned 3D priors could address these limitations, but training them at scale is challenging due to memory and runtime constraints. This work proposes SPUR-iG, a fully 3D deep unrolled subspace reconstruction framework that makes large-scale non-Cartesian 3D MRF computationally feasible.
Methods: SPUR-iG leverages implicit GROG-based data consistency (DC), which grids non-Cartesian k-space using a learned family of kernels, enabling efficient FFT-based DC with minimal artifacts. To enable scalable 3D unrolled training, we introduce a staged training strategy that keeps computation tractable while progressively improving reconstruction quality. We evaluate the method on a large in vivo dataset, as well as on cross-vendor out-of-distribution data.
@@
Conclusion: SPUR-iG introduces a fully 3D unrolled reconstruction framework for MRF that improves both reconstruction speed and accuracy, making high-resolution accelerated 3D MRF more practical for research and clinical use.
@@ 1 Introduction
Magnetic Resonance Fingerprinting (MRF) ma2013magnetic is a qMRI technique that encodes tissue parameters through a sequence of varying flip angles (FA), repetition times (TR), and potentially other acquisition settings across $N_{\text{TR}}$ time points. At each TR, an image is acquired using an incoherency-promoting trajectory, producing artifacts that are spatially and temporally incoherent and thus reduce interference with tissue-specific signal evolutions. After reconstructing the time series of images, voxel-wise signal evolutions are matched to a dictionary of Bloch-simulated signals, linking tissue parameters to their expected dynamics and enabling parameter estimation.
@@
Subspace reconstruction zhao2018improved, asslander2018low, liang2007spatiotemporal approximates the signal dictionary with a low-rank subspace of size $k\ll N_{\text{TR}}$, thereby lowering dimensionality and implicitly regularizing the reconstruction. The resulting subspace coefficient maps provide a compact representation of the image’s signal evolutions across TRs, which can be used to extract quantitative $T_{1}$ and $T_{2}$ maps. Beyond traditional parameter mapping, these coefficient maps capture rich signal information that facilitates additional downstream tasks, such as multi-compartment modeling nagtegaal2020fast, and high-quality clinical contrast synthesis yurt2024unlocking, wang2023high. This motivates focusing on accurate subspace coefficient reconstruction that can serve as input to many downstream applications.
For highly accelerated MRF, additional regularization beyond the subspace model is required. Locally Low Rank (LLR) tamir2017t2, lima2019sparsity is widely used, enforcing low-rank structure within local patches, but it requires computationally expensive singular value decompositions (SVD) cai2010singular. At very high acceleration rates, such as acquisitions targeting 1 mm isotropic whole-brain mapping in under 2 minutes, LLR alone is insufficient to preserve reconstruction quality.
Broadly, achieving high-fidelity 3D MRF at these resolutions and acceleration factors is limited by two primary bottlenecks: (1) the reliance on handcrafted priors that lack the expressivity to resolve fine details under high undersampling, and (2) the computationally expensive NUFFTs and regularization operators, which substantially raise reconstruction time and memory usage at high resolutions.
@@
To overcome limitations of the LLR prior, state-of-the-art unrolled methods such as SOTIP zou2025improved incorporate learned priors via deep learning, yielding substantial gains in image quality. To accelerate reconstruction, SOTIP employs NUFFTs with reduced kernel sizes and oversampling factors, which introduce small artifacts that the denoiser must learn to remove. Additionally, although data consistency (DC) is performed in 3D, denoising is applied slice-by-slice in 2D to reduce computational cost.
While such design choices are often necessary for computational efficiency, slice-wise denoising limits the ability to fully exploit the inherently 3D structure of MRF data, which exhibits volumetric artifacts and spatial correlations. More generally, several works have explored strategies to enable fully 3D learned reconstruction under memory constraints. Memory-efficient unrolled frameworks based on reversible networks and gradient checkpointing have demonstrated 3D structural MRI reconstruction with substantially reduced memory footprints, at the expense of increased computation and additional numerical considerations in reversible DC and model layers kellman2020memory, wang2021memory. Other approaches reduce memory by reformulating the reconstruction. Efficient Cartesian methods exploit separable structure to process smaller reformatted volumes deng2021efficient, while greedy layer-wise training schemes perform backpropagation after each unroll iteration ozturkler2022gleam, reducing memory by roughly a factor equal to the number of unroll steps, while keeping the overall computational cost approximately unchanged. However, because intermediate iterates are supervised independently rather than through the final unrolled objective, the optimization target may not fully align with the final reconstruction quality. Spatiotemporal architectures have further extended learned reconstruction to Cartesian dynamic imaging by factoring spatial and temporal processing to improve memory utilization kustner2020cinenet.
While these approaches represent important advances, the challenges are further compounded in high-resolution MRF, where the reconstruction problem is inherently 3D, spatiotemporal, and non-Cartesian, limiting the direct application of the above methods.
To address these issues, we propose SPUR-iG (Staged Pretraining for Unrolled Reconstruction with iGROG), a fully 3D deep unrolled reconstruction framework aggarwal2018modl, adler2018learned, hammernik2018learning, sriram2020end. Our approach combines efficient iGROG-based DC with a learned 3D prior and a staged training strategy that enables large-scale unrolled training. We first pretrain a standalone denoiser with extensive data augmentation, then train with a greedy per-unroll-iteration loss to reduce memory demands ozturkler2022gleam, and finally fine-tune the full unrolled model using gradient checkpointing chen2016training. To further improve robustness, the denoiser takes as input both the image and the unroll iteration index, enabling it to amortize performance across varying artifact and noise levels. Evaluated on 45 in vivo MRF acquisitions with retrospective undersampling spanning acquisition times from 2 minutes down to 30 seconds, and including an example from an out-of-distribution scanner vendor, our framework achieves whole-brain 1 mm isotropic reconstructions in under 15 s in all cases, while surpassing existing methods in quality.
@@ 2 Methods
To enable fast and accurate MRF reconstruction, we build on the standard subspace formulation and address its limitations of suboptimal priors and long reconstruction times. We propose a fully 3D unrolled framework that integrates a learned prior with an efficient DC formulation, along with a staged training strategy that makes large-scale training feasible within a reasonable compute and time budget. We begin by reviewing MRF acquisition and the subspace model, followed by unrolled reconstruction, and then present the details of our proposed method and evaluation pipeline. An overview of our approach is illustrated in Figure 1.
Figure 1:  Illustration of inference (top) and training (bottom) pipelines of SPUR-iG. During inference, iGROG grids the non-Cartesian data onto a Cartesian grid using an implicit kernel representation learned from the calibration region. The gridded k-space is then passed to the unrolled reconstruction, initialized with ${\bf\cal A}^{H}{\bf b}$. Reconstruction proceeds for NN unroll iterations, each alternating between a DC update with a learnable step size $\mu_{i}$ and a step-conditioned denoiser. Training is performed in three stages: (1) a step-conditioned UNet is pretrained on inputs with varying artifact levels, using a heuristic mapping between input type and unroll step to promote robustness across noise/artifact conditions; (2) the pretrained UNet is used to initialize a memory-efficient GLEAM-style training, where losses are computed and backpropagated after each unroll iteration, detaching subsequent steps; and (3) full unrolled training, initialized from stage (2), fine-tunes the model under the exact inference setting, with gradient checkpointing (GC) across DC and denoising steps to reduce memory consumption.
@@ 2.1 MRF Subspace Reconstruction
MRF consists of a sequence of $N_{\text{TR}}$ acquisitions with varying excitation parameters (e.g., FAs), producing a time series of images ${\bf x}\in\mathbb{C}^{N_{\text{TR}}\times N_{v}}$, where $N_{v}$ denotes the number of voxels in the volume. Reconstruction typically follows two stages: (1) recovery of the time series $\hat{{\bf x}}$ from undersampled k-space data, and (2) voxel-wise parameter estimation via dictionary matching, where the temporal signal evolution is compared to a precomputed dictionary, and quantitative maps are obtained by selecting the entry with maximum cosine similarity.
Directly reconstructing the full time series $\bf x$} is computationally intensive and ill-posed due to its high dimensionality and the heavy undersampling used in MRF. This can be mitigated by exploiting temporal redundancy and projecting the data onto a low-dimensional subspace ${\bm{\Phi}}\in\mathbb{C}^{N_{\text{TR}}\times k}$ ($k\ll N_{\text{TR}}$) spanned by the top $k$ singular vectors of the MRF dictionary. The time series is then approximated as ${\bf x}\approx{\bm{\Phi}}{\bm{\alpha}}$, where ${\bm{\alpha}}\in\mathbb{C}^{k\times N_{v}}$ are the subspace coefficient maps. We define the forward operator acting on $\bm{\alpha}$} as ${\bf\cal A}={\bf\cal F}_{u}{\bf S}{\bm{\Phi}}$, with $\bf S$} the coil sensitivities and ${\bf\cal F}_{u}$ the undersampled Fourier operator. Since ${\bf\cal F}_{u}{\bf S}$ acts on the spatial domain and $\bm{\Phi}$} on the temporal domain, these operators commute, allowing the costly ${\bf\cal F}_{u}{\bf S}$ to be applied directly in the reduced $k$-dimensional space tamir2017t2.
@@ 2.2 Unrolled Reconstruction
Equation (1) is commonly solved using accelerated proximal algorithms such as FISTA beck2009fast, which alternate between DC updates and proximal steps. However, at very high acceleration rates, LLR regularization becomes insufficient to constrain the undersampled problem, and in addition, its proximal operator requires repeated SVDs of local patches, making it computationally expensive.
Unrolled reconstruction aggarwal2018modl, adler2018learned, hammernik2018learning, sriram2020end addresses these issues by replacing the proximal operator with a learned denoiser, yielding the update scheme
@@ 2.3.1 Efficient Data Consistency via iGROG
The main cost in (2) arises from evaluating ${\bf\cal A}^{H}{\bf\cal A}(\cdot)$, which relies on NUFFTs. To accelerate this step, we use iGROG, which learns an INR of the gridding kernel across multi-channel coils. Using calibration data, the INR is trained to map $n$ offset vectors ${\bf d}_{1},\ldots,{\bf d}_{n}$, corresponding to displacements between non-Cartesian samples and a target Cartesian grid point, to a kernel $G_{{\bm{\psi}}^{*}}({\bf d}_{1},\ldots,{\bf d}_{n})$. For each target grid location, nearby source samples are selected, the kernel is evaluated at their relative displacements, and the resulting kernel weights are used to compute the target k-space value. This enables GRAPPA-like griswold2002generalized interpolation and, through the use of an INR, generalizes across the diverse orientations encountered in the trajectory. This allows us to first grid the non-Cartesian data onto a Cartesian grid, after which NUFFTs can be replaced by efficient FFTs. The upper left part of Figure 1 illustrates this process.
The approach provides flexibility to trade off accuracy and noise amplification against speed. By using an oversampled grid, interpolation quality improves (since the interpolation distances are smaller), but reconstruction time increases. In our experiments, we used a $\times 1.5$ oversampling ratio with 5 k-space data points per coil channel, achieving high-quality reconstructions while retaining substantial speed-ups. Across all acquisitions considered in this work, INR training completed in under 25 s and can be performed during MRF acquisition immediately after calibration data are acquired. Once trained, gridding of the dataset is likewise efficient, completing in under 3 s and yielding data on a Cartesian grid for subsequent FFT-based DC updates.
@@ 2.3.2 Efficient Training of a 3D Unrolled Network
To enable training of a 3D unrolled network in the challenging MRF reconstruction setting, we adopt a progressive three-stage strategy, illustrated in the lower part of Figure 1. This strategy consists of the following stages:
@@
    Denoiser pretraining: We begin by pretraining a standalone, iteration-conditioned 3D UNet on a diverse set of inputs, where the model is trained to recover clean images from noisy, undersampled inputs. To obtain a robust initialization for subsequent stages, we expose the network to a wide range of artifact and noise levels at its input, including zero-filled (${\bf\cal A}^{H}{\bf b}$) and LLR-based reconstructions from scans with different acquisition durations (30 s, 1 min, 2 min, and 6 min, denoted ${\bm{\alpha}}_{\text{LLR}}^{\text{30s}},\ldots,{\bm{\alpha}}_{\text{LLR}}^{\text{6min}}$). Since the UNet is iteration-conditioned, but this stage does not involve actual unrolling and therefore has no natural iteration index, we assign heuristic pseudo-iteration indices to each input type. Specifically, ${\bf\cal A}^{H}{\bf b}$ is paired with index 1, corresponding to the initialization of unrolled reconstruction, while reconstructions with increasing acquisition durations are mapped to higher indices (e.g., ${\bm{\alpha}}_{\text{LLR}}^{\text{30s}}\to 2$). Because this stage does not involve expensive unrolling or DC updates, we can efficiently apply extensive data augmentation, including random rotations, translations, and scaling. At the end of this stage, the network provides a well-initialized denoising prior that generalizes across a range of artifact levels and incorporates a notion of the unroll iteration index.
@@
    Greedy unrolled training (GLEAM ozturkler2022gleam): Starting from the pretrained weights, we train the unrolled network using a greedy scheme, in which the loss is backpropagated through only a single unroll iteration at a time, after which the computation graph is detached. This avoids storing the full unrolled graph, which would otherwise exceed typical GPU memory. Since the final output of the unrolled process is of primary interest, we assign geometrically increasing loss weights across iterations, emphasizing later steps while allowing greater flexibility in earlier ones. Importantly, this stage resolves the input distribution mismatch introduced during prior pretraining: while the denoiser from Stage 1 is trained on reconstructed images with different artifacts levels, during unrolled inference its true inputs are the intermediate outputs of preceding iterations. By the end of this stage, we obtain a fully unrolled model that achieves high reconstruction quality, with faster training convergence and improved performance due to the informed initialization.
@@
Collectively, these stages enable efficient large-scale training by progressively mitigating the limitations of each preceding stage, while circumventing the need to train a fully 3D unrolled model from scratch, which would otherwise be prohibitively compute expensive and slow. This process ultimately yields a high-quality, fully 3D unrolled reconstruction model. We refer to the resulting framework as SPUR-iG, introduced in Section 1. Code will be made publicly available upon publication.